Payrolls for the 3 highest paid baseball teams (the Yankees, Red Sox, and Angels) for 2006 totaled 418 million. If the team payroll for the Yankees is 75 million dollars higher than the Red Sox and the Red sox is 17 million dollars higher than the Angels, then what was the 2006 payroll for each team?
My work:
Yankees= 75+R (R stands for Red Sox)
Red Sox= 17+A (A stands for Angels)
Angels= X
X+17+X+75+X=418
X=108.666666667 or 108.67
Angels= 108.67
Red Sox= 125.67
Yankees= 200.67
When added together it comes out to: 435.01
I am close to the 418, I just can't figure out where i messed up. I assume in my equation.
Thanks

Answer by stanbon(75887)

Payrolls for the 3 highest paid baseball teams (the Yankees, Red Sox, and Angels) for 2006 totaled 418 million. If the team payroll for
the Yankees is 75 million dollars higher than the Red Sox and the
Red sox is 17 million dollars higher than the Angels, then what was the 2006 payroll for each team?
My work:
Yankees= 75+R = 75+17+A = 92+A (R stands for Red Sox)
Red Sox= 17+A (A stands for Angels)
Angels= X
------
Equation:
A + 17+A + 92+A = 418
3A + 109 = 418
3A = 309
A = 103
----------
etc.
Cheers,
Stan H.